15:30 Hall went to Controlled Access for LMS network troubleshooting. Back to Beam Permit at 1630.
16:35 Mark Ito reports attempts to get experts in to work on the no-beam problem: they'll be in tomorrow. We can't get beam from Hall C as no one wants to take away beam from G0 (then have the beam go down), and no one wants to definitively say we can't have it, either. Jay has been convinced to come in on Saturday evening and look at the situation: he'll give a more definitive answer after he knows what's going on.
16:40 Crew Chief called & says Hall C is down & they're willing to try and steer us something.
17:35 Jay Bennet called & says he's sending us pulsed beam.
17:45 Jay called - tagger won't go & stay on. Tagger expert & Mark Ito were called.
18:00 Krister Bruhwel calls in - relayed to MCC. He and the Engineering on-call guy are coming in.
18:25 Krister came in. He took controlled access to the hall in order to fix the tagger magnet problems.
20:25 Krister calls with a request for MCC to ramp the tagger magnet - he's done it locally and wants MCC to try. MCC reports a timeout and Crister is given the MCC phone number for direct interaction.
20:45 Denny and Krister traced the problem to Q11 on the Mosfet Gating and Monitoring Board. The board that had the problem was the second from the left on the bottom side. After replacing this mosfet the Tagger ramped to field and is now under MCC's control.
21:00 Mark Ito contacted & he says go to Beam Permit and ask MCC for 2-5 nA beam into the tagger dump. MCC responds that we're working parasiticly with Hall C and they'll do what they can.
21:15 MCC says they'll try to give us 0.5 microamp of pulsed beam.
21:16 MCC says they'll try to send 2-5 nanoAmps of CW beam.
21:19 Bright spot appears in the Tagger Dump. Intermittently, and the gamma profiler never shows a peak.
21:24 MCC calls & says they're having problems. Hall C is continuously running so diagnostics are limited: he'll do the best he can, but it'll probably be only a few nanoamps.
21:34 MCC informs they've got the beam into the tagger magnet, so they're going to CW beam. Bright spot appears in the Tagger Dump but gamma profiler still shows no peak.
21:39 Spot in Tagger Dump vanishes.
21:41 MCC calls and says the best they can give us is 170 - tagger megnet's limited to 180 or so. BPM position screen bpm1nA_BPM1Hz.adl reports (right colum 2H01) XPOS and YPOS as Out of Range (ignore this comment - Liping says MCC reported that one as dead).
22:10 Folks are here, Mark Ito consulted, alarms silenced, 5-pass beam at 4.7 GeV established. They're still centering the beam.
22:30 Beam centered. Continuing to follow the run plan.
22:57 Problems with Harp Tagger scan. Hall B beamline expert contacted & is coming in.
24:00 Shift end. David Lawrence brought in, beam stops, various problems with software under analysis.